OpenStudy (anonymous):

What is the value of x? x +5= -6x -16

OpenStudy (anonymous):

-7x=-21 x=3

OpenStudy (michele_laino):

sorry, x=-3

OpenStudy (uri):

x +5= -6x -16 x+6x=-16-5 Solve further? @rcwitt00

OpenStudy (anonymous):

7x=-21 x=-3 thanks @Michele_Laino

OpenStudy (michele_laino):

thanks!

OpenStudy (uri):

It's not a race. Try explaining the person instead. @Michele_Laino @koook

OpenStudy (michele_laino):

@uri I'm sorry, you are right! sorry again and thanks for your advice

OpenStudy (uri):

yw. @rcwitt00 Would you LIKE to respond? We are trying to help you out.